ian has 6 nature books.for each nature book, he has 2 science books.for every science book, he has 2 animal books.how many animal books does he have?
Answer by Theo(13342)

i believe it will be 6 * 2 * 2 = 24

2 science books for each nature book gives him 6 * 2 = 12 science books.

2 animal books for each science book give him 12 * 2 = 24 animal books.